dots sells 250 t-shirts for $2.00 and shorts for $4.00. In april total sales were $600.00. How many t-shirts and shorts did dots sell.

Answer by edjones(7569)

250*2=$500. worth of teeshirts.
600-500=$100. worth of shorts.
100/4=25 shorts sold
.
Ed